# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ 4.2 Slab-experts for hinge loss
$\leq\left(1+2\eta\gamma\right)\left(\sum\limits_{i=1}^{T}\frac{(\gamma-y_{i}B_{{\theta^{\star}},\alpha,\gamma}(x_{i}))_{+}}{\gamma}+\frac{9d}{\eta\gamma}\log(eT/d)\right)$
$\leq(1+2\eta\gamma)\left((1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}\langle x,\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\eta\gamma}+\frac{3d}{\eta\gamma}\log\left(\pi T\right)\Bigg{)}+\frac{9d}{\eta\gamma}\log(eT)\right)$
$\leq(1+2\eta\gamma)^{2}\left(\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},{\theta^{\star}}\rangle)_{+}}{\gamma}\right)+\frac{12d}{\eta\gamma}(1+2\eta\gamma)^{2}\log(\pi T)+\frac{1}{\eta\gamma}(1+2\eta\gamma)^{2}.$
@@ 5 Computational Aspects
In this section, we shortly discuss the computational properties of our algorithms in the problems considered so far, namely, sparse linear regression, logistic regression, and hinge loss regression.
@@ 5.1 Computation for sparse linear regression
In this section, we discuss the computational properties of our methods in Theorem 2.2 and Proposition A.1. For the sparse prior exponential weights algorithm considered in Theorem 2.2, we compare it to the SeqSEW algorithm proposed by Gerchinovitz (2013). The SeqSEW algorithm considers a similar sparse prior as in (7), but the exponential weights algorithm is applied to the loss generated with clipped predictions, i.e., $\ell_{\theta}(x_{t},y_{t})=\left(y_{t}-\operatorname{clip_{m}}\left(\langle\theta,x_{t}\rangle\right)\right)^{2}$. This clipped loss is no longer convex, thus sampling from the posterior remains unclear. In contrast, our exponential weights run on the loss of the unclipped prediction; thus, the convexity still holds. Concretely, we would like to compute the quantity $\widehat{f}_{t}(x_{t})$ as defined in (8), which is a ratio between two normalizing constants for log-concave distributions. To begin, we first assume that $A_{t}=\sum_{i=1}^{t}x_{i}x_{i}^{\top}$ is full rank. Then, it is straightforward to check that both functions are
$\frac{1}{2m^{2}}(m-\langle x,\theta\rangle)^{2}+\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}+4\sum_{j=1}^{d}\log\left(1+|\theta^{(j)}|\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau\right)~{}~{}\text{and}$
@@
$\frac{1}{2m^{2}}(-m-\langle x,\theta\rangle)^{2}+\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}+4\sum_{j=1}^{d}\log\left(1+|\theta^{(j)}|\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau\right)$
are $\lambda_{\min}(A_{t})/(2m^{2})$-strongly convex and $\left(\lambda_{\max}(A_{t})/(2m^{2})+4\max_{j}\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}/\tau^{2}\right)$-smooth. Let $\xi_{t}=\lambda_{\min}(A_{t})/\left(\lambda_{\max}(A_{t})+8m^{2}\max_{j}\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}/\tau^{2}\right)$ be the condition number. Thus, by the recent result of Ge, Lee, and Lu (2020, Theorem 1.1), normalizing constants for both log-concave densities can be estimated with a multiplicative factor of $1\pm\varepsilon$ within $\widetilde{O}\left(\frac{d^{4/3}\xi_{t}+d^{7/6}\xi_{t}^{7/6}}{\varepsilon^{2}}\right)$ number of steps, where $\widetilde{O}(f)$ for any growth function $f$ stands for $O(f)\cdot\log^{O(1)}(f)$. It suffices to estimate $\widehat{f}(x_{t})$ up to an additive error of $O(m/T)$. Thus, by choosing $\varepsilon=O(1/T)$, we can estimate $\widehat{f}_{t}(x_{t})$ up to accuracy of $m/2\cdot(\log(1+\varepsilon)-\log(1-\varepsilon))\leq 2m\varepsilon$. This, in turn, shows that at each step, estimating $\widehat{f}(x_{t})$ requires only
$\widetilde{O}\left(\left(d^{4/3}\xi_{t}+d^{7/6}\xi_{t}^{7/6}\right)T^{2}\right)$
number of steps. In general, we can add a small amount of quadratic regularization to the prior distribution to ensure strong convexity. Consider the prior distribution
@@
$\widetilde{\mu}(\theta)=c_{0}\exp\left(-\lambda\sum\nolimits_{t=1}^{T}\langle x_{t},\theta\rangle^{2}\right)\prod\limits_{j=1}^{d}\frac{3\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau}{2\left(1+|\theta^{(j)}|\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau\right)^{4}},$ (39)
where $\lambda$ is left to be specified. Since $\exp\left(-\lambda\sum\nolimits_{t=1}^{T}\langle x_{t},\theta\rangle^{2}\right)\leq 1$, we know that $c_{0}\geq 1$. For this prior, we consider the estimator
[…]
This estimator results in the following regret bound.
Proposition 5.1.
Assume that the sequence $\{(x_{t},y_{t})\}_{t=1}^{T}$, where $x_{t}\in\mathbb{R}^{d}$ and $y_{t}\in\mathbb{R}$ for all $1\leq t\leq T$, is such that $\max_{t}|y_{t}|\leq m$, $\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$ is full rank, and that the smallest scaled singular value condition is satisfied with constant $\kappa_{s}$. Then, for any $s$-sparse $\theta^{\star}\in\mathbb{R}^{d}$, the following regret bound holds for the estimator (40) with $\tau=m\sqrt{\frac{s}{d}}$ and $\lambda=\frac{1}{2m^{2}}\left(\frac{s}{T}\wedge 1\right)$,
@@
$\sum\limits_{t=1}^{T}(y_{t}-\widetilde{f}_{t}(x_{t}))^{2}\leq\sum\limits_{t=1}^{T}(y_{t}-\langle x_{t},\theta^{\star}\rangle)^{2}+sm^{2}\left(6+8\log\left(1+\frac{2}{\kappa_{s}}\sqrt{\frac{dT}{s}}\right)\right)$
We remark that for the computationally efficient version of the predictor, we assume the full covariance matrix $\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$ is of full rank. While this may seem contradictory to classical sparsity results, where the number of observations is much smaller than the dimensionality of the data, it is less of an issue in the sequential setup. Here, we consider a long sequence of rounds $T$, such that $T\geq d$, while the comparator vector $\theta^{\star}$ is sparse. One motivation for this is that, unlike classical methods, which focus on recovering parameters with limited data, sequential methods benefit from the accumulation of observations, where the ongoing data accumulation corresponds to a very long horizon $T$, possibly much larger than $d$.
Proof of Proposition 5.1 The proof is identical to the proof of Theorem 2.2 with the only difference in the estimation of $\mathcal{KL}(\mu^{\star}\|\widetilde{\mu})$ instead of $\mathcal{KL}(\mu^{\star}\|\mu)$. Concretely, we have
$\mathcal{KL}(\mu^{\star}\|\widetilde{\mu})$ $=\int_{\mathbb{R}^{d}}\log\frac{\mu^{\star}(\theta)}{\widetilde{\mu}(\theta)}d\mu^{\star}(\theta)$
$=\mathcal{KL}(\mu^{\star}\|\mu)+\int_{\mathbb{R}^{d}}\log\frac{\mu(\theta)}{\widetilde{\mu}(\theta)}d\mu^{\star}(\theta)$
$=\mathcal{KL}(\mu^{\star}\|\mu)+\lambda\int_{\mathbb{R}^{d}}\sum\nolimits_{t=1}^{T}\langle x_{t},\theta\rangle^{2}d\mu^{\star}(\theta)-\log c_{0}$
$\leq\mathcal{KL}(\mu^{\star}\|\mu)+\lambda\int_{\mathbb{R}^{d}}\sum\nolimits_{t=1}^{T}\langle x_{t},\theta\rangle^{2}d\mu^{\star}(\theta),$
where the last inequality is by the fact that $c_{0}\geq 1$. Furthermore, by the definition of $\mu^{\star}$, we have
$\int_{\mathbb{R}^{d}}\sum\nolimits_{t=1}^{T}\langle x_{t},\theta\rangle^{2}d\mu^{\star}(\theta)$ $=\int_{\mathbb{R}^{d}}\sum\nolimits_{t=1}^{T}\langle x_{t},\theta+{\theta^{\star}}\rangle^{2}d\mu(\theta)$
$=\sum\nolimits_{t=1}^{T}\langle x_{t},{\theta^{\star}}\rangle^{2}+\int_{\mathbb{R}^{d}}\sum\nolimits_{t=1}^{T}\langle x_{t},\theta\rangle^{2}d\mu(\theta)$
$=\sum\nolimits_{t=1}^{T}\langle x_{t},{\theta^{\star}}\rangle^{2}+\tau^{2}d$
$\leq 4m^{2}T+\tau^{2}d.$
Thus, we obtain
@@
$\sum\limits_{t=1}^{T}(y_{t}-\widetilde{f}_{t}(x_{t}))^{2}\leq\sum\limits_{t=1}^{T}(y_{t}-\langle x_{t},\theta^{\star}\rangle)^{2}+sm^{2}\left(1+2m^{2}\lambda+8\log\left(1+\frac{2}{\kappa_{s}}\sqrt{\frac{dT}{s}}\right)\right)+8\lambda m^{4}T.$
Plugging in $\lambda$, we obtain the result. ∎
For our sparse prior of (43) considered in Proposition A.1, the computation aspect is similar to Rigollet and Tsybakov (2012), which is discussed in their Section 7.
@@ 5.2 Computation for logistic regression
Motivated by the high computational complexity of the algorithm proposed by Foster et al. (2018), the computational problem of online logistic regression has been further studied in (Jézéquel et al., 2020, 2021). However, since their algorithms are designed for the online setup, the lower bound in Theorem 3.1 shows that they cannot achieve the regret upper bound achieved by our Algorithm 1. On the other hand, the Normalized Maximum Likelihood (NML) algorithm is known to achieve exact minimax regret (Wu et al., 2022), but its computation requires exhausting all possible label sequences $y_{1},\ldots,y_{T}\in\{\pm 1\}^{T}$, which scales as $2^{T}$. Consequently, without additional algorithmic insights, the NML algorithm cannot be computed in polynomial time. Methods based on $\varepsilon$-nets typically work for the problems we consider but suffer from the lack of known implementable algorithms constructing the $\varepsilon$-nets (see Appendix D for detailed discussion).
@@ 5.2.1 Computing slab-expert prediction
For any $\theta\in\mathbb{R}^{d}$, to produce our prediction for the slab-expert $B_{\theta,\alpha}$ as defined in Definition 3.1 at time step $t$ on a point $x_{t}$ inside a slab $K_{\theta}\subset\{x_{i}\}_{i\in[T]}$ , we need to estimate probability assignments defined as $B_{\theta,\alpha}(x_{t},1)=\int\sigma(-\langle x_{t},\theta\rangle)\pi_{t}(d\theta)$ where the density is
$\pi_{t}(\theta)\propto\exp\left(\sum\limits_{x\in K_{t-1}}\log(\sigma(y_{x}\langle x,\theta\rangle))-\alpha\theta^{\top}\left(\sum\nolimits_{x\in K}xx^{\top}\right)\theta\right),$
@@ 5.2.2 Aggregating the slab-experts
Definition 5.1 (Generalized slab-expert).
Let $S_{1}=(K_{1,+},K_{1,-})$ and $S_{2}=(K_{2,+},K_{2,-})$ be any two elements in the set of the half-space separations $\mathcal{S}$. If $K_{1,-}\cap K_{2,+}=\emptyset$, then denote by $A_{S_{1},S_{2}}$ the generalized slab-expert that runs exponential weights on set $K=K_{1,+}\cap K_{2,-}=\{x_{t}\}_{t\in[T]}\setminus(K_{1,-}\cup K_{2,+})$, predicts $-1$ on $K_{1,-}$, and predicts $1$ on $K_{2,+}$. Denote the set of all such generalized slab-experts by $\mathcal{A}^{+}$.
It is clear that the slab experts are such experts as well. Thus, $\mathcal{A}\subseteq\mathcal{A}^{+}$. Since the number of halfspaces is upper bounded by $|\mathcal{S}|\leq(eT/d)^{d}$, we have $|\mathcal{A}^{+}|\leq|\mathcal{S}|^{2}\leq(eT/d)^{2d}$. The generalized slab-experts are constructed through all possible halfspaces. This reduces the task of aggregating slab-experts to the simpler problem of enumerating/sampling halfspaces. Indeed, in the following, we show that it is possible to enumerate all the halfspaces with the help of the SVM algorithm. Moreover, it is possible to use rejection sampling where each halfspace will be chosen with probability at least $1/(2T)^{d+1}$ and with acceptance at least $1/2$. This, together with the Metropolis-Hastings (MH) algorithm for evaluating exponential weighted averaging, constitutes an implementable algorithm.
@@
To enumerate the generalized slab-experts, we only need to enumerate all the possible halfspaces. Fortunately, the hard-margin SVM for halfspaces in $\mathbb{R}^{d}$ with $n$ input labeled data points can output at most $d+1$ labeled data points that capture all the labeling information (specified in the proof of Lemma 5.1) with computational cost $O(\mathrm{poly}(d,n))$ (Boser et al., 1992). Utilizing this fact, we can enumerate all the halfspaces for any given covariate set in polynomial time, as shown by the following lemma.
Lemma 5.1.
For any set ${K}$ of $T>0$ covariates, the halfspaces restricted to ${K}$ defined as $\mathrm{HalfSpace}_{K}=\{h_{\theta,b}|_{K}:h_{\theta,b}=\operatorname{sign}(\langle x,\theta\rangle-b),\theta\in\mathbb{R}^{d},b\in\mathbb{R}\}$ can be enumerated using at most $O((2T)^{d})$ calls to the $\mathrm{SVM}$ algorithm.
@@
We first enumerate all the subsets of ${K}$ with at most $d+1$ elements, i.e., ${\mathcal{K}}_{d+1}=\{{K}^{\prime}\subseteq{K}:|{K}^{\prime}|\leq d+1\}$. For each of the subset ${K}^{\prime}$ in ${\mathcal{K}}_{d+1}$, we consider the set of all possible labelings $\mathcal{D}({K}^{\prime})=\{\{\pm 1\}^{{K}^{\prime}}:{K}^{\prime}\in{\mathcal{K}}_{d+1}\}$. For any of the labeling $\mathcal{D}=\{(x,y_{x})\}_{x\in{K}^{\prime}}\in\mathcal{D}({K}^{\prime})$, we can apply the $\mathrm{SVM}$ algorithm to obtain $(\widehat{\mathcal{D}},\widehat{h})=\mathrm{SVM}(\mathcal{D})$. With this process, we can define the following set of linear separators
$\mathcal{H}_{{K},\mathrm{SVM}}=\left\{\widehat{h}|_{{K}}\;\middle|\;\begin{aligned} (\widehat{\mathcal{D}},\widehat{h})=\mathrm{SVM}(\mathcal{D})&,~{}~{}\mathcal{D}\in\mathcal{D}({K}^{\prime}),\\ \mathcal{D}({K}^{\prime})=\{\{\pm 1\}^{{K}^{\prime}}:{K}^{\prime}\in{\mathcal{K}}_{d+1}\}&,~{}~{}{\mathcal{K}}_{d+1}=\{{K}^{\prime}\subseteq{K}:|{K}^{\prime}|\leq d+1\}\end{aligned}\right\}.$
We claim that $\mathrm{HalfSpace}_{K}=\mathcal{H}_{{K},\mathrm{SVM}}$. For any hypothesis $h|_{K}\in\mathrm{HalfSpace}_{K}$, let $(\widehat{\mathcal{D}},\widehat{h})=\mathrm{SVM}(\{x,h(x)\}_{x\in{K}})$. By the property of $\mathrm{SVM}$, we have $|\widehat{\mathcal{D}}|\leq d+1$, $h|_{K}=\widehat{h}|_{K}$ and $(\widehat{\mathcal{D}},\widehat{h})=\mathrm{SVM}(\widehat{\mathcal{D}})$.
@@
Proposition 5.2.
Algorithm 1, when using generalized slab-experts $\mathcal{A}^{+}$ instead of the slab-experts $\mathcal{A}$, can be implemented with a computational cost that scales with at most $O(\mathrm{poly}(T^{d}))$ and achieves a regret upper bound of $O(d\log T)$.
Proof of Proposition 5.2 Applying Lemma 5.1, we know that we can enumerate the halfspaces on given covariate set with computational cost $O(\mathrm{poly}(T^{d}))$. Once the halfspaces are enumerated, the generalized slab-experts can be constructed by the combinations of all halfspaces. Then Algorithm 1 with generalized slab-experts $\mathcal{A}^{+}$ aggregates each round with sampling cost at most $O(\mathrm{poly}(T,d))$ as shown in Section 5.2.1 and computational cost at most $O(\mathrm{poly}(T^{d}))$ (since there are $\mathrm{poly}(T^{d})$ generalized slab-experts). In all, Algorithm 1, when using generalized slab-experts $\mathcal{A}^{+}$ instead of the slab-experts $\mathcal{A}$, can be implemented with a computational cost that scales as at most $O(\mathrm{poly}(T^{d}))$. Moreover, recall that the number of generalized slab-experts is upper bounded by $(eT/d)^{2d}$. Thus, the uniform prior on the generalized slab-experts puts at least $\Omega(\mathrm{poly}(1/T^{d}))$ probability on the optimal slab-expert $B_{{\theta^{\star}},\alpha}$. Therefore, the regret can be upper bounded by $O(d\log T)$. The claim follows. ∎

### v3

@@ 4.2 Slab-experts for hinge loss
$\leq\left(1+2\eta\gamma\right)\left(\sum\limits_{i=1}^{T}\frac{(\gamma-y_{i}B_{{\theta^{\star}},\alpha,\gamma}(x_{i}))_{+}}{\gamma}+\frac{4d}{\eta\gamma}\log(eT/d)\right)$
$\leq(1+2\eta\gamma)\left((1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}\langle x,\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\eta\gamma}+\frac{3d}{\eta\gamma}\log\left(\pi T\right)\Bigg{)}+\frac{4d}{\eta\gamma}\log(eT)\right)$
$\leq(1+2\eta\gamma)^{2}\left(\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},{\theta^{\star}}\rangle)_{+}}{\gamma}\right)+\frac{7d}{\eta\gamma}(1+2\eta\gamma)^{2}\log(\pi T)+\frac{1}{\eta\gamma}(1+2\eta\gamma)^{2}.$
@@ 5 Computational Aspects
In this section, we briefly discuss the computational properties of our algorithms in the problems considered so far, namely, sparse linear regression, logistic regression, and hinge loss regression.
@@ 5.1 Computation for sparse linear regression
In this section, we discuss the computational properties of our methods in Theorem 3 and Proposition 34. Concretely, we would like to compute the quantity $\widehat{f}_{t}(x_{t})$ as defined in (8). We provide three possible approaches for this computation. First, set
${L}_{m}(\theta)$ $=\frac{1}{2m^{2}}(m-\langle x_{t},\theta\rangle)^{2}+\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2},$
${L}_{-m}(\theta)$ $=\frac{1}{2m^{2}}(-m-\langle x_{t},\theta\rangle)^{2}+\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}.$
The first approach is to notice that computing $\widehat{f}_{t}(x_{t})$ is equivalent to computing
@@
$\frac{\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{m}(\theta)\right)\mu(\theta)d\theta}{\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{-m}(\theta)\right)\mu(\theta)d\theta}=\int_{\mathbb{R}^{d}}\frac{\exp\left(-V(\theta)\right)}{\int_{\mathbb{R}^{d}}\exp\left(-V(\theta^{\prime})\right)d\theta^{\prime}}\cdot\exp\left(\frac{2}{m}\langle x_{t},\theta\rangle\right)d\theta,$
where
$V(\theta)={L}_{-m}(\theta)+4\sum_{j=1}^{d}\log\left(1+|\theta^{(j)}|\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau\right).$
Thus, we can sample from the distribution with density proportional to $\exp\left(-V(\theta)\right)$, following almost verbatim the approach in Dalalyan and Tsybakov (2012b, Section 4), who utilized the Langevin Monte Carlo method [^3] Note that, in our setting, it remains unclear whether the Langevin Monte Carlo method admits a worst-case polynomial-time implementation. This limitation also applies to the results of Dalalyan and Tsybakov (2012b).. Using these samples, we can evaluate the expectation of $\exp\left(\frac{2}{m}\langle x_{t},\theta\rangle\right)$ under that distribution.
The second approach is to notice that
@@
$\frac{\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{m}(\theta)\right)\mu(\theta)d\theta}{\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{-m}(\theta)\right)\mu(\theta)d\theta}=\frac{\mathbb{E}_{\theta\sim\xi_{1}}\left[\prod_{j=1}^{d}\left({1+|\theta^{(j)}|\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau}\right)^{-4}\right]}{\mathbb{E}_{\theta\sim\xi_{2}}\left[\prod_{j=1}^{d}\left({1+|\theta^{(j)}|\cdot\sqrt{\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2}}/\tau}\right)^{-4}\right]},$
where
@@
$\xi_{1}(\theta)$ $\propto\exp\left(-\frac{1}{2m^{2}}(m-\langle x_{t},\theta\rangle)^{2}-\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}\right),$
@@
$\xi_{2}(\theta)$ $\propto\exp\left(-\frac{1}{2m^{2}}(-m-\langle x_{t},\theta\rangle)^{2}-\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}\right)$
are both Gaussian distributions. If sufficiently many Gaussian samples from $\xi_{1}$ and $\xi_{2}$ are obtained to approximate both expectations up to multiplicative constants close to one, the value of $\widehat{f}_{t}(x_{t})$ can then be approximated within the desired precision level.
Finally, we can analyze Laplace’s approximation, which is computationally efficient. Specifically, we can compute two least squares solutions, $\widehat{\theta}_{m}=\arg\min_{\theta\in\mathbb{R}^{d}}{L}_{m}(\theta)$ and $\widehat{\theta}_{-m}=\arg\min_{\theta\in\mathbb{R}^{d}}{L}_{-m}(\theta)$. When the asymptotic error of the Laplace’s method is small, i.e.,
$\begin{cases}\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{m}(\theta)\right)\mu(\theta)\,d\theta\approx\exp\left(-{L}_{m}(\widehat{\theta}_{m})\right)\mu(\widehat{\theta}_{m}),\\ \int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{-m}(\theta)\right)\mu(\theta)\,d\theta\approx\exp\left(-{L}_{-m}(\widehat{\theta}_{-m})\right)\mu(\widehat{\theta}_{-m}).\end{cases}$
Using these solutions, we can express the following approximate identity:
$\frac{\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{m}(\theta)\right)\mu(\theta)\,d\theta}{\int\limits_{\mathbb{R}^{d}}\exp\left(-{L}_{-m}(\theta)\right)\mu(\theta)\,d\theta}\approx\frac{\exp\left(-{L}_{m}(\widehat{\theta}_{m})\right)\mu(\widehat{\theta}_{m})}{\exp\left(-{L}_{-m}(\widehat{\theta}_{-m})\right)\mu(\widehat{\theta}_{-m})}.$
We leave a rigorous analysis of the last two approaches as an interesting direction for future research.
@@ 5.2 Computation for logistic regression
Motivated by the high computational complexity of the algorithm proposed by Foster et al. (2018), the computational problem of online logistic regression has been further studied in (Jézéquel et al., 2020, 2021). However, since their algorithms are designed for the online setup, the lower bound in Theorem 9 shows that they cannot achieve the regret upper bound achieved by our Algorithm 1. On the other hand, the Normalized Maximum Likelihood (NML) algorithm is known to achieve exact minimax regret (Wu et al., 2022), but its computation requires exhausting all possible label sequences $y_{1},\ldots,y_{T}\in\{\pm 1\}^{T}$, which scales as $2^{T}$. Consequently, without additional algorithmic insights, the NML algorithm cannot be computed in polynomial time. Methods based on $\varepsilon$-nets typically work for the problems we consider but suffer from the lack of known implementable algorithms constructing the $\varepsilon$-nets (see Appendix D for detailed discussion).
@@ 5.2.1 Computing slab-expert prediction
For any $\theta\in\mathbb{R}^{d}$, to produce our prediction for the slab-expert $B_{\theta,\alpha}$ as defined in Definition 11 at time step $t$ on a point $x_{t}$ inside a slab $K_{\theta}\subset\{x_{i}\}_{i\in[T]}$ , we need to estimate probability assignments defined as $B_{\theta,\alpha}(x_{t},1)=\int\sigma(\langle x_{t},\theta\rangle)\pi_{t}(d\theta)$ where the density is
$\pi_{t}(\theta)\propto\exp\left(\sum\limits_{x\in K_{t-1}}\log(\sigma(y_{x}\langle x,\theta\rangle))-\alpha\theta^{\top}\left(\sum\nolimits_{x\in K_{\theta}}xx^{\top}\right)\theta\right),$
@@ 5.2.2 Aggregating the slab-experts
Definition 22 (Generalized slab-expert).
Let $S_{1}=(K_{1,+},K_{1,-})$ and $S_{2}=(K_{2,+},K_{2,-})$ be any two elements in the set of the half-space separations $\mathcal{S}$. If $K_{1,-}\cap K_{2,+}=\emptyset$, then denote by $A_{S_{1},S_{2}}$ the generalized slab-expert that runs exponential weights on set $K=K_{1,+}\cap K_{2,-}=\{x_{t}\}_{t\in[T]}\setminus(K_{1,-}\cup K_{2,+})$, predicts 1-1 on $K_{1,-}$, and predicts $1$ on $K_{2,+}$. Denote the set of all such generalized slab-experts by $\mathcal{A}^{+}$.
It is clear that the slab experts are such experts as well. Thus, $\mathcal{A}\subseteq\mathcal{A}^{+}$. Since the number of half-spaces is upper bounded by $|\mathcal{S}|\leq(eT/d)^{2d}$, we have $|\mathcal{A}^{+}|\leq|\mathcal{S}|^{2}\leq(eT/d)^{4d}$. The generalized slab-experts are constructed through all possible halfspaces. This reduces the task of aggregating slab-experts to the simpler problem of enumerating/sampling halfspaces. Indeed, in the following, we show that it is possible to enumerate all the halfspaces with the help of the SVM algorithm. Moreover, it is possible to use rejection sampling where each halfspace will be chosen with probability at least $1/(2T)^{d+1}$ and with acceptance at least $1/2$. This, together with the Metropolis-Hastings (MH) algorithm for evaluating exponential weighted averaging, constitutes an implementable algorithm.
@@
In order to enumerate the generalized slab-experts, we only need to enumerate all the possible halfspaces. Fortunately, the hard-margin SVM for halfspaces in $\mathbb{R}^{d}$ with $n$ input labeled data points can output at most $d+1$ labeled data points that capture all the labeling information (specified in the proof of Lemma 23) with computational cost $O(\mathrm{poly}(d,n))$ (Boser et al., 1992). Utilizing this fact, we can enumerate all the halfspaces for any given covariate set in polynomial time, as shown by the following lemma.
Lemma 23.
For any set K{K} of $T>0$ covariates, the halfspaces restricted to K{K} defined as $\mathrm{HalfSpace}_{K}=\{h_{\theta,b}|_{K}:h_{\theta,b}=\operatorname{sign}(\langle x,\theta\rangle-b),\theta\in\mathbb{R}^{d},b\in\mathbb{R}\}$ can be enumerated using at most $O((2T)^{d})$ calls to the $\mathrm{SVM}$ algorithm.
@@
We first enumerate all the subsets of K{K} with at most $d+1$ elements, i.e., ${\mathcal{K}}_{d+1}=\{{K}^{\prime}\subseteq{K}:|{K}^{\prime}|\leq d+1\}$. For each of the subset ${K}^{\prime}$ in ${\mathcal{K}}_{d+1}$, we consider the set of all possible labelings $\mathcal{D}({K}^{\prime})=\{\{\pm 1\}^{{K}^{\prime}}:{K}^{\prime}\in{\mathcal{K}}_{d+1}\}$. For any of the labeling $\mathcal{D}=\{(x,y_{x})\}_{x\in{K}^{\prime}}\in\mathcal{D}({K}^{\prime})$, we can apply the $\mathrm{SVM}$ algorithm to obtain $(\widehat{\mathcal{D}},\widehat{h})=\mathrm{SVM}(\mathcal{D})$. With this process, we can define the following set of linear separators
$\mathcal{H}_{{K},\mathrm{SVM}}=\left\{\widehat{h}|_{{K}}\;\middle|\;\begin{aligned} (\widehat{\mathcal{D}},\widehat{h})=\mathrm{SVM}(\mathcal{D})&,\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{D}\in\mathcal{D}({K}^{\prime}),\\ \mathcal{D}({K}^{\prime})=\{\{\pm 1\}^{{K}^{\prime}}:{K}^{\prime}\in{\mathcal{K}}_{d+1}\}&,\leavevmode\nobreak\ \leavevmode\nobreak\ {\mathcal{K}}_{d+1}=\{{K}^{\prime}\subseteq{K}:|{K}^{\prime}|\leq d+1\}\end{aligned}\right\}.$
[…]
@@
Proposition 24.
Algorithm 1, when using generalized slab-experts $\mathcal{A}^{+}$ instead of the slab-experts $\mathcal{A}$, can be implemented with a computational cost $O(\mathrm{poly}(T^{d}))$ and achieves a regret upper bound of $O(d\log T)$.
Proof of Proposition 24 Applying Lemma 23, we know that we can enumerate the halfspaces on given covariate set with computational cost $O(\mathrm{poly}(T^{d}))$. Once the halfspaces are enumerated, the generalized slab-experts can be constructed by the combinations of all halfspaces. Then Algorithm 1 with generalized slab-experts $\mathcal{A}^{+}$ aggregates each round with sampling cost at most $O(\mathrm{poly}(T,d))$ as shown in Section 5.2.1 and computational cost at most $O(\mathrm{poly}(T^{d}))$ (since there are $\mathrm{poly}(T^{d})$ generalized slab-experts). In all, Algorithm 1, when using generalized slab-experts $\mathcal{A}^{+}$ instead of the slab-experts $\mathcal{A}$, can be implemented with a computational cost that scales as at most $O(\mathrm{poly}(T^{d}))$. Moreover, recall that the number of generalized slab-experts is upper bounded by $(eT/d)^{4d}$. Thus, the uniform prior on the generalized slab-experts puts at least $\Omega(\mathrm{poly}(1/T^{d}))$ probability on the optimal slab-expert $B_{{\theta^{\star}},\alpha}$. Therefore, the regret can be upper bounded by $O(d\log T)$. The claim follows. ∎